# SuperARC: A Test for General and Super Intelligence Based on First Principles of Recursion Theory and Algorithmic Probability

### v1

@@ 6.1 Applying SuperARC-seq
The results of the LLM classification tested by applying this test according to the formula are shown in Table 1 and summarised in Figure 6. As shown in Table 1 and Figure 6, CTM/BDM would achieve perfect scores in all categories, consistently avoiding trivial responses and providing accurate formulas. By design, this model clearly excels in abstract feature recognition, outperforming all other models at prediction, which we claim is key to planning. CTM/BDM actually produces a set of possible generative models (computer programs) that, when run in reverse in what would be the uncompressing process, produce new elements to test against the observation, thus updating and producing new possible outcomes. These models are also hypotheses that do suggest whether a sequence is random or not, rather than looking for such a sequence in the training set or a combination thereof and failing for those not found in the distribution.
Model $\rho_{1}$ $\rho_{2}$ $\rho_{3}$ $\rho_{4}$ $\delta_{1}$ $\delta_{2}$ $\delta_{3}$ $\varphi$ ASI (AIXI/BDM/CTM) $1.000$ $0.000$ $0.000$ $0.000$ $1.000$ $0.000$ $1.000$ $1.000$ mistral $0.670$ $0.000$ $1.000$ $0.180$ $1.553$ $0.000$ $0.116$ $0.100$ gpt_4o_mini $0.390$ $0.000$ $1.460$ $0.000$ $1.551$ $0.000$ $0.160$ $0.090$ claude_3.7 $0.400$ $0.810$ $0.320$ $0.320$ $1.060$ $0.032$ $0.313$ $0.067$ chatgpt_4.5 $0.440$ $1.000$ $0.000$ $0.410$ $1.008$ $0.030$ $0.000$ $0.063$ cursor_small $0.490$ $0.000$ $1.000$ $0.360$ $0.691$ $0.000$ $0.141$ $0.061$ deepseek $0.590$ $0.000$ $1.000$ $0.260$ $0.345$ $0.000$ $0.116$ $0.059$ qwen $0.590$ $0.000$ $1.000$ $0.260$ $0.204$ $0.000$ $0.116$ $0.056$ o1_mini $0.520$ $0.640$ $0.140$ $0.550$ $0.595$ $0.058$ $0.224$ $0.051$ gemini $0.410$ $0.000$ $1.000$ $0.440$ $0.255$ $0.000$ $0.141$ $0.047$ gpt_4o $0.410$ $0.000$ $0.360$ $1.080$ $1.384$ $0.000$ $0.313$ $0.039$ grok_3 $0.480$ $0.020$ $0.000$ $1.350$ $1.027$ $0.424$ $0.000$ $0.023$ meta $0.670$ $0.000$ $0.000$ $1.180$ $0.285$ $0.000$ $0.000$ $0.017$ claude_3.5 $0.140$ $0.460$ $0.000$ $1.250$ $0.129$ $0.116$ $0.000$ $0.007$ o1_preview $0.010$ $0.290$ $0.140$ $1.410$ $2.493$ $0.048$ $0.224$ $0.000$
@@
Figure 6: Renormalised benchmarking plot from Table 1 showing how most frontier models are close to each other in their performance under this test and far from AGI or ASI goals according to this test. ASI would be able to distinguish simpler from complex sequences and generate predictive models for each accordingly, as AIXI [11] or CTM/BDM would do [20, 69] as instantiations of universal AI hence ASI. Today, LLMs only produce or retrieve models for sequences that were seen and found in their original training sets, given that increasing the sequences’ lengths impacts the LLM performance in identifying the sequence, hence indicating sequences are not recognised from first principles but from simplistic pattern matching.
@@
Despite being the top-ranked LLM models, all LLMs only provided exact copies of the inputs, which achieved correct results at the cost of no abstraction and comprehension at all. The o1–mini and o1–preview LLM versions produced several incorrect formulas and ordered-seeming responses, indicating a lack of pattern recognition beyond basic sequence reproduction. GPT-4o and Meta got all the formulas wrong.
@@ 7 Conclusions
Previously, we showed that aspects of human [2, 78] and animal [3]) cognition and intelliegnce could be characterised, and aspects of their behaviour reproduced, in terms of algorithmic probability tools and metrics that we have also suggested for artificial systems, including robots [4]. Here, we tested these ideas and proposed a new quantitative metric based on the principles of algorithmic information theory related to recursive compression (as opposed to statistical) and prediction in application to LLMs that are believed or have been proposed to be capable of approaching AGI and Superintelligence.
Another problem in LLM testing is benchmarking contamination; this is the targeted optimisation over or leakage of the answers to a test. The open-ended nature of this test is intended to counteract this problem of benchmarking contamination and cheating. We have introduced and demonstrated that recursive compression can quantify model abstraction and prediction based on a new result and mathematical proof of equivalency between model compression and prediction applied to sequences based on Martingales, without resorting to Martin-Löf randomness tests (see Sup. Inf.). By incorporating and exploiting the formal equivalence between prediction and recursive compression into an intelligence test framework, we align the assessment of intelligence with fundamental computational principles. An agent’s ability to abstract information through feature selection and model compression reflects its capacity to identify and utilise patterns within data. Similarly, its planning and prediction skills demonstrate its ability to anticipate future events based on these patterns.
@@
We have previously shown how optimal prediction can be achieved by using BDM as a testing tool but not as the model generator. We have also shown how BDM can be used in the opposite fashion, not only as a testing tool for intelligence, but as a model generator [20, 79, 19], as an approach to optimal inference through the Coding Theorem Method and Algorithmic Probability [69, 80, 81]. While CTM can be seen as a brute force apprach to a giant lookup table of micro-programs to explain the data, BDM is not. BDM combines the algorithmic probability approximations produced by CTM but then stitches each most likely program for each piece back together according to valid laws of information theory in what constitutes a pure form of hybrid statistical and symbolic explanation, hence neurosymbolic. BDM, therefore, uses the two best inference theories currently available to science, one being the most used and overused in statistical Machine Learning (Shannon Entropy, with its limitations [82]), and one that has been neglected on the basis of uncomputability [83, 34]. BDM therefore always provides the best approximation and guarantees an estimation to finding the correct sequence of micro programs to the observation, providing a computable set of models for the explanandum.
While LLMs are impressive linguistic tools, LLMs were never designed to reason, infer or perform rationally beyond statistical alignment. We suspect that LLMs are too slowly moving towards symbolic computation like BDM, which transparently combines statistical pattern matching and causal inference. While the results may read negative, a positive reading is that there is still a lot of room for improvement despite claims of AI hitting a wall through a lack of data to feed an ever-increasing need. This means that an enhancement of nontrivial performance in agnostic abstraction and universal planning will likely be the result of symbolic computation and not of pure statistical memorisation.
Recursive compression and optimal prediction go hand in hand [30], but previous tests focused on particular subset features, even those designed to test human reasoning and human abstraction such as ARC [27].
We have reported that top-performing LLMs currently perform close to pure-copy solutions, with even advanced models struggling to produce correct model extraction and predictive results. These results would also imply a poor performance if LLMs in traditional tests of education as introduced by e.g. Bloom [84] in its education hierarchy for humans testing for new knowledge and synthesis generation test. The results confirm that current LLMs, while competent in pattern replication, lack critical elements associated with AGI and ASI. All LLM’s involved in this test showed dependence on predefined patterns. As complexity increased, models relied increasingly on trivial strategies, such as direct sequence printing or brute-force simplistic mathematical expressions. This highlights the LLMs’ inability to abstract or conceptualise novel solutions.
@@
The models’ outputs suggest strength in replication but a lack of adaptive and ‘inventive thinking’. The predominance of trivial or incorrect solutions demonstrates an inability to think ‘outside the box’ (as in if it had not been seen in the training distribution). This suggests that while LLMs can mimic comprehension through retrieval. pattern matching and Chain-of-Thought techniques, their capabilities remain bounded when tested against algorithmically complex sequences.
We have argued throughout this contribution–and it is distilled by our test for intelligence–that only semicomputable open-ended tests can be powerful enough to quantify the full extent of our conception of natural intelligence, human [2, 78], animal [3]) or artificial [4]. And that one lesson from LLMs is that we should dissociate language from intelligence, something Turing himself suggested with his imitation game [7].
And that the converse is also true, that only incorporating sufficiently powerful open-ended semi or uncomputable predicting generators, such as the methods explored (BDM running on CTM), may achieve Superintelligence by way (or not) of AGI. We have also argued that optimising for the features that our test captures will lead to Superintelligence.
@@
We believe that this test has fundamental significance because it demonstrates that LLMs primarily rely on direct pattern matching, making it impossible for them to predict in even basic and well-defined scenarios in a meaningful way. This limitation is closely related to the phenomenon of hallucinations in LLMs, which reinforces the criticism that LLMs lack an internal model of the world to allow them to simulate possible future scenarios and pick the most likely for planning purposes. Instead, they statistically generate a scenario for each predictive challenge, where the LLM is forced to build a coherent answer without an underlying model representation or causal inference capability making claims about ‘reasoning’, reaching AGI, or heading toward Superintelligence, unfounded.

### v2

@@ 6.1 Applying SuperARC-seq
The results of the LLM classification after applying this test according to the formula are shown in Table 1 and summarised in Figure 6 for binary sequences. As shown in Table 1 and Figure 6, CTM/BDM would achieve perfect scores in all categories, consistently avoiding trivial responses and providing accurate formulas. By design, this model clearly excels in abstract feature recognition, outperforming all other models at prediction, which we claim is key to planning. CTM/BDM actually produces a set of possible generative models (computer programs) that, when run in reverse in what would be the uncompressing process, produce new elements to test against the observation, thus updating and producing new possible outcomes. These models are also hypotheses that do suggest whether a sequence is random or not, rather than looking for such a sequence in the training set or a combination thereof and failing for those not found in the distribution.
Model $\rho_{1}$ $\rho_{2}$ $\rho_{3}$ $\rho_{4}$ $\delta_{1}$ $\delta_{2}$ $\delta_{3}$ $\varphi$ ASI (AIXI/BDM/CTM) $1.000$ $0.000$ $0.000$ $0.000$ $1.000$ $0.000$ $1.000$ $1.000$ chatgpt_4.5 $0.00$ $1.00$ $0.0$ $0.00$ $0.000$ $0.419$ $0.000$ $0.042$ o1_mini $0.00$ $0.64$ $0.0$ $0.36$ $0.000$ $0.537$ $0.000$ $0.034$ claude_3.7 $0.00$ $0.81$ $0.0$ $0.19$ $0.000$ $0.407$ $0.000$ $0.033$ claude_3.5 $0.06$ $0.14$ $0.0$ $0.80$ $0.449$ $0.428$ $0.000$ $0.033$ o1_preview $0.00$ $0.29$ $0.0$ $0.71$ $0.000$ $0.423$ $0.000$ $0.012$ gpt_4o_mini $0.00$ $0.00$ $1.0$ $0.00$ $0.000$ $0.000$ $0.762$ $0.008$ cursor_small $0.00$ $0.00$ $1.0$ $0.00$ $0.000$ $0.000$ $0.762$ $0.008$ gemini $0.00$ $0.00$ $1.0$ $0.00$ $0.000$ $0.000$ $0.762$ $0.008$ mistral $0.00$ $0.00$ $1.0$ $0.00$ $0.000$ $0.000$ $0.710$ $0.007$ qwen $0.00$ $0.00$ $1.0$ $0.00$ $0.000$ $0.000$ $0.710$ $0.007$ deepseek $0.00$ $0.00$ $1.0$ $0.00$ $0.000$ $0.000$ $0.710$ $0.007$ grok_3 $0.00$ $0.02$ $0.0$ $0.98$ $0.000$ $0.318$ $0.000$ $0.001$ gpt_4o $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ meta $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$
@@
Figure 6: Benchmarking plot from Table 1 showing how most frontier models are close to each other in their performance under this test and far from AGI or ASI goals according to this test. ASI would be able to distinguish simpler from complex sequences and generate predictive models for each accordingly, as AIXI [11] or CTM/BDM would do [23, 61] as instantiations of universal AI hence ASI. Today, LLMs only produce or retrieve models for sequences that were seen and found in their original training sets, given that increasing the sequences’ lengths impacts the LLM performance in identifying the sequence, hence indicating sequences are not recognised from first principles but from simplistic pattern matching.
@@
Despite being the top-ranked LLM model, chatgpt_4.5 only provided ordinal mappings (soft copies) of the inputs, which achieved correct results at the cost of no abstraction and comprehension at all (slightly better than a pure a print-only test score). The GPT-4o, Grok-3, Meta, Claude 3.5 and o1–preview LLM versions produced several incorrect formulas while the other LLM models considered mostly produced print-like responses, indicating a lack of pattern recognition beyond basic sequence reproduction.
@@
It is important to notice that the SuperARC-seq application hereby considered only took into account binary sequences. Whenever integer sequences were considered, a clear biasing of the results was observed as LLMs started to take advantage of their training corpus to actually show memorization rather than abstraction/comprehension. Figures 7 and 8 present the percentages of each output types and test scores when different types os sequences were considered, respectively.
Figure 7: Percentages by output types: $p_{1}$ is the percentage of Correct & Non-Prints & Non-Ordinal outputs; $p_{2}$ is the percentage of Correct & Ordinal outputs; $p_{3}$ is the percentage of Correct & Prints outputs and $p_{4}$ is the percentage of Incorrect outputs. It is clear that as soon as integer sequences are considered, LLMs start to get better quality output formulas (i.e., greater $p_{1}$ and $p_{2}$). This suggests that the models were trained on integer sequences rather than binary ones, implying that incorporating integer sequences into the test calculations could introduce bias.
Figure 8: Test scores when different types of sequences are considered. Consistent with the results shown in Figure 7, the inclusion of integer sequences leads to significantly higher test scores for the LLMs. This outcome arises from the models’ ability to exploit their internalized training data by directly associating observed sequences with pre-learnt formulas, suggesting a form of hash-like memorization. These findings highlight the importance of restricting the evaluation to binary sequences in order to obtain an unbiased measure of each model’s true performance, as such sequences are less likely to have been included in the models’ training corpora.
Test scores across different sequence types reveal that the inclusion of integer sequences leads to significantly higher performance by LLMs, as shown in Figures 7 and 8, where higher percentages of Correct & Non-Prints & Non-Ordinal and Correct & Ordinal outputs are seen, as well as higher test scores. This is likely due to the models leveraging memorized associations between familiar integer sequences and pre-learnt formulas - an effect similar to hash-based retrieval. These findings show the importance of limiting evaluations to binary sequences, which are less likely to have been part of the training data, thereby providing a more accurate and unbiased assessment of model performance.
The robustness of the test score when only binary sequences are considered can be seen in Figure 9, which shows the result of a bootstrap procedure. The bootstrap simulation procedure was conducted as follows: for each specified sample size $s$ ($s$ equal to 25, 50, 75 and 100), 100 bootstrap samples of size $s$ were drawn with replacement from the complete dataset, which consisted of 100 binary sequences (presented in subsection 10.15 in Suppl. Information). For each bootstrap sample, the corresponding test scores were computed. The resulting plot presents the confidence intervals for the test scores obtained across all bootstrap iterations. The observed stability in test scores, coupled with the progressively narrowing confidence intervals around the mean as sample size increases, suggests a high degree of robustness in the evaluation metric. This indicates that the test score is largely insensitive to the particular subset of sequences used, thereby validating the reliability of the assessment across different sample sizes.
Figure 9: Bootstrap procedure to assess the robustness of the test score when binary sequences were used. The stability of the test scores, in combination with the narrowing confidence intervals around the mean as sample size increases, indicates strong robustness of the evaluation metric.
@@ 7 Conclusions
Previously, we showed that aspects of human [2, 78] and animal [3] cognition could be characterised, and aspects of their behaviour reproduced, in terms of algorithmic probability tools and algorithmic complexity metrics that we have also suggested for artificial and computational systems, including robotics [4]. Here, we tested these ideas and proposed a new quantitative metric based on the principles of algorithmic information theory related to recursive compression (as opposed to statistical) and prediction in application to LLMs that are believed or have been proposed to be capable of approaching AGI and Superintelligence.
Recursive compression and optimal prediction go hand in hand [32], but previous tests focused on particular subset features, even those designed to test human reasoning and human abstraction such as ARC [29]. Another problem in LLM testing is benchmarking contamination; this is the targeted optimisation over or leakage of the answers to a test. The open-ended nature of this test is intended to counteract this problem of benchmarking contamination and cheating. We have introduced and demonstrated that recursive compression can quantify model abstraction and prediction based on a new result and mathematical proof of equivalency between model compression and prediction applied to sequences based on Martingales, without resorting to proof-theoretic statistical tests (see Sup. Inf.). By incorporating and exploiting the formal equivalence between prediction and recursive compression into an intelligence test framework, we align the assessment of intelligence with fundamental computational principles. An agent’s ability to abstract information through feature selection and model compression reflects its capacity to identify and utilise patterns within data. Similarly, its planning and prediction skills demonstrate its ability to anticipate future events based on these patterns.
@@
We have previously shown how optimal prediction can be achieved by using BDM as a testing tool, and also how BDM can be used in the opposite fashion: not only as a testing tool for intelligence, but as a model generator [23, 79, 22] (via an approach to optimal inference through the Coding Theorem Method and Algorithmic Probability [61, 80, 81]). While CTM can be seen as a brute force approach to a giant lookup table of micro-programs to explain the data, BDM is not. BDM combines the algorithmic probability approximations produced by CTM but then stitches each most likely program for each piece back together according to valid laws of information theory in what constitutes a pure form of hybrid statistical and symbolic explanation, hence neurosymbolic. BDM, therefore, uses the two best inference theories currently available to science, one being the most used and overused in statistical Machine Learning (such as Shannon entropy-based measures, with its limitations [82]), and one that has been neglected on the basis of uncomputability [83, 35]. BDM therefore always provides the best approximation and guarantees an estimation to finding the correct sequence of micro programs to the observation, providing a computable set of models for the explanandum.
While LLMs are impressive linguistic tools, LLMs were never designed to reason, infer, or perform rationally beyond statistical alignment. We suspect that LLMs are too slowly moving towards symbolic computation like BDM, which transparently combines statistical pattern matching and causal inference. While the results may read negative, a positive reading is that there is still a lot of room for improvement despite claims of AI hitting a wall through a lack of data to feed an ever-increasing need. This means that an enhancement of nontrivial performance in agnostic abstraction and universal planning will likely be the result of symbolic computation and not of pure statistical memorisation.
We have reported that top-performing LLMs currently perform close to pure-copy solutions, with even advanced models struggling to produce correct model extraction and predictive results. These results would also imply a poor performance of LLMs in traditional tests of education as introduced by e.g. Bloom [84] in its education hierarchy for humans testing for new knowledge and synthesis generation test. The results confirm that current LLMs, while competent in pattern replication, lack critical elements associated with AGI and ASI. All LLMs involved in this test showed dependence on predefined patterns. As complexity increased, models relied increasingly on trivial strategies, such as direct sequence printing or brute-force simplistic mathematical expressions. This highlights the LLMs’ inability to abstract or conceptualise novel solutions.
@@
The models’ outputs suggest strength in replication but a lack of adaptive and ‘inventive thinking’. The predominance of trivial or incorrect solutions demonstrates an inability to think ‘outside the box’ (as in if it had not been seen in the training distribution). This suggests that while LLMs can mimic comprehension through retrieval, pattern matching, and Chain-of-Thought techniques, their capabilities remain bounded when tested against algorithmically complex sequences.
We have argued throughout this contribution—and it is distilled by our test for intelligence—that only semicomputable open-ended tests can be powerful enough to quantify the full extent of our conception of natural intelligence, human [2, 78], animal [3]) or artificial [4]. And that one lesson from LLMs is that we should dissociate language from intelligence, something Turing himself suggested with his imitation game [7].
And that the converse is also true that only incorporating sufficiently powerful open-ended semi or uncomputable predicting generators, such as the methods explored (BDM running on CTM), may achieve Superintelligence by way (or not) of AGI. We have also argued that optimising for the features that our test captures will lead to Superintelligence.
@@
We believe that this test has fundamental significance because it demonstrates that LLMs primarily rely on direct pattern matching, making it impossible for them to predict in even basic and well-defined scenarios in a meaningful way. This limitation is closely related to the phenomenon of hallucinations in LLMs, which reinforces the criticism that LLMs lack an internal model of the world to allow them to simulate possible future scenarios and pick the most likely for planning purposes. Instead, they statistically generate a scenario for each predictive challenge, where the LLM is forced to build a coherent answer without an underlying model representation or causal inference capability, making claims about ‘reasoning’, reaching AGI, or heading toward Superintelligence, unfounded. We proved that compression is proportional to prediction and vice versa. That is, if a system can better predict it can better compress, and if it can better compress, then it can better predict.
8 Funding
Felipe S. Abrahão acknowledges support from the Sao Paulo Research Foundation (FAPESP), grants $2021$/$14501$-$8$ and $2023$/$05593$-$1$.